### v1

@@ 1 Introduction
From human hairs to cables for electronic devices, elastic strands can be found everywhere in our daily lives, and modeling such thread-like materials is of critical importance in various domains, including computer graphics, mechanical engineering, and medical applications. The Discrete Elastic Rods (DER) formulation [Bergou2008, Bergou:2010:DVT:1778765.1778853] is one of the most popular approaches to accurately and efficiently simulate such elastic strands and has demonstrated its effectiveness in a wide range of applications involving both forward simulations [Chentanez2009needle, DiVerdi2010, AUDOLY201318, Kaufman2014, Schweickart2017, Fei2019, Lesser2022, Daviet2023hair] and inverse problems [Perez2015rod, Zehnder2016curve, Perez2017surfaces, Schumacher2018sheet, Malomo2018flexmaps, Xu2018wire, Panetta2019, Neveu2022curve, Ren2022umbrella, chen2024differentiablediscreteelasticrods].
When modeling and designing elastic strands, a well-known problem is that such strands will immediately sag due to gravity when the simulation starts [Twigg:2011:OSS:2019406.2019437, Hsu2023sag], which ruins the specific strand shapes (e.g., hairstyles) that were either carefully designed by users, technical artists, and designers [Iben2019, Liu2022gravity] or captured from real human hair or wigs [Luo2013hair]. This issue occurs because the simplest way to model the strands is to focus solely on the geometry, neglecting the role of forces or dynamics; that is, the strand’s initial shape is treated as its rest shape [Kim2022deformables]. While one can largely mitigate this sagging effect by significantly (and artificially) increasing the material stiffness, this approach makes the elastic strands excessively stiff and sacrifices their natural dynamic response, thus negatively impacting the visual result.
To address this sagging problem in the DER framework, we propose a rest shape optimization approach that optimizes the rest shape parameters (rest length, rest curvature, and rest twist) of the input strands such that the input shape is preserved at static equilibrium under simulation (i.e., the strand shape deformed due to simulation coincides with the input shape). To find optimal rest shape parameters, we formulate a minimization problem based on the kinetic energy. Given the greater number of degrees of freedom (DOFs) for the rest shape parameters compared to the DOFs of DER strands, we incorporate a regularizer to avoid creating underdetermined systems. In addition, we impose box constraints on the rest shape parameters to prevent significant changes from their original values, ensuring the stability of the system. We solve the resulting constrained nonlinear minimization problem via the Gauss-Newton (GN) algorithm augmented with penalty methods and line search. Fig.  Author Guidelines for EUROGRAPHICS Proceedings Manuscripts demonstrates the efficacy of our rest shape optimization.
@@ 2.1 Elastic Rod Simulation
Due to the one-dimensional structure of thin elastic strands, it is typically difficult to accurately simulate these materials with general finite element methods (FEM) designed for volumetric structures [Kim2022deformables], and specialized approaches have been investigated. Given that common strand-like materials exhibit higher stiffness in stretching modes compared to bending and twisting, early work assumed inextensibility via e.g., constraints based on Lagrange multipliers [Goldenthal2007cloth] with linear-time factorization and triangular solves [Baraff1996linear], a tridiagonal matrix formulation [Han2013tridiagonal], and geometric stiffness [Tournier2015, Andrews2017stiffness], reduced multi-body solvers [Hadap2006strands, Featherstone2016], and position-based dynamics (PBD) [Muller:2007:PBD:1234415.1234575, Bender2017survey]. To support the characteristic dynamics of elastic strands, the traditional mass-spring framework [Liu:2013:FSM:2508363.2508406] was extended with altitude springs [Selle2008mass], bending springs [Iben2013hair], torsion energy [Shi2023curls], and ghost rest shapes [Herrera2024]. To simulate one-dimensional strands along with other deformables in a unified way, supporting torsional effects is critical, and therefore discretization elements that can capture twists (e.g., oriented particles [Muller2011oriented] and elastons [Martin2010unified]) have also been presented. For highly constrained strand-like systems, further specialized approaches have been proposed with, e.g., Eulerian nodes [Sueda2011strands, Sanchez2020rods], long range constraints [Muller2017rigid], and unilateral distance constraints [Muller2018cable].
Since the introduction of the Cosserat theory into graphics [Pai2002cosserat], it has been adapted to dynamical systems of strands [Spillmann2007rod] and network structures [Spillmann2009nets], combined with PBD using ghost particles [Umetani2014rod], and further extended to support edge rotations via quaternions in the PBD framework [Kugelstadt2016rod]. The work of Kugelstadt and Schömer [Kugelstadt2016rod] has been extended to remove the ambiguity in the Darboux vector [Hsu2023sag], to handle stiff problems via compliant constraints while utilizing the tree structures [Deul2018rod], to preserve the volume of elastic strands [Angles2019rods], and to achieve consistent dynamics [Soler2018rod] using projective dynamics [Bouaziz:2014:PDF:2601097.2601116]. Elastic rod simulation based on the Cosserat theory [Spillmann2007rod] has also been extended to support rh-adaptivity [Wen2020rod] and a compact representation of the DOFs [Zhao2022rod]. A curvature-based reduced representation has also been presented by Bertails et al. [Bertails2006hair] and extended to generate smoother dynamics via clothoids [Casati2013clothoids].
While various approaches have been presented for simulating elastic strands, DER [Bergou2008, Bergou:2010:DVT:1778765.1778853], which discretizes an elastic strand with vertex positions and edge angles in maximal coordinates, is one of the most popular approaches due to its simplicity, efficiency, and flexibility, as employed in a wide range of applications. As such, we use DER as our elastic strand simulator.
@@ 2.2.1 Nonlinear Force Solve
Solving Newton’s second law of motion for (typically nonlinear) forces and zero acceleration is equivalent to achieving zero net force and thus static equilibrium. Hadap [Hadap2006strands] proposed using inverse dynamics (which computes forces from acceleration) [Featherstone2016] to solve the nonlinear equation on the forces, and thus enable sag-free simulation. While inverse dynamics works for reduced multibody systems, in general, it is not clear how to apply it for systems represented in maximal coordinates.
The nonlinear equations on the forces can be solved, e.g., with Newton-type optimizers [Wang:2015:DCM:2809654.2766911, Mukherjee2018shape], although Iben [Iben2019] aimed to cancel gravity using spring forces via local analytical solutions within the mass-spring systems [Iben2013hair]. However, in practice, it is necessary to consider a global system to fully cancel gravity. A similar technique was also employed for sag-free mass-spring systems by introducing additional artificial springs (which can potentially introduce unnatural forces) [Herrera2024].
Chen et al. [Chen2014asymptotic] employed an Asymptotic Numerical Method (ANM), which is a homotopy approach to solving nonlinear equations, for inverse elastic shape design with incompressible neo-Hookean materials. They demonstrated orders of magnitude faster convergence than one of the Newton-type optimizers, Levenberg–Marquardt algorithm (LMA) [NoceWrig06]. Later, Jia [Jia2021SANM] extended ANM to support more general material types and also demonstrated its efficiency for forward simulation compared to optimization-based integrators [Martin:2011:EEM:2010324.1964967, Gast2015] combined with Newton’s method [NoceWrig06] (although ANM can yield solutions that are local maxima, unlike those generated by the optimization-based integrators). ANM was also employed to investigate the equilibrium of elastic rods [LAZARUS20131712]. While ANM can be more efficient than the Newton-type optimizers [NoceWrig06], it is not clear how to support box constraints, which are indispensable to constrain rest shape parameters within physically valid ranges and avoid introducing stability issues.
@@ 2.2.2 Constrained Minimization
Derouet-Jourdan et al. [Derouet-Jourdan2010] presented a method for achieving static equilibrium of elastic strands, simulated with the curvature-based discretization [Bertails2006hair], in 2D. They formulated an objective consisting of curvatures and material parameters and then optimized these parameters via linear solves on the curvature and stiffening/lightening of materials. Derouet-Jourdan et al. [Derouet-Jourdan:2013:IDH:2508363.2508398] also achieved static equilibrium of curvature-based hair strands [Bertails2006hair] in 3D under frictional contacts. Along the same lines, inverse approaches for curvature-based elastic strands have been further investigated [bertailsdescoubes:hal-01827887, Charrondiere2020, Hafner2021curves, Hafner2023rod].
Twigg and Kačić-Alesić proposed optimizing rest shape parameters via force-norm minimization to achieve static equilibrium of mass-spring systems [Twigg:2011:OSS:2019406.2019437]. Our work is most closely related to theirs, so we clarify the differences in detail in Sec. 2.2.4. Skouras et al. [Skouras2012design] presented an optimization approach that treats the nonlinear force equations as hard constraints and solves the optimization via the augmented Lagrangian method [NoceWrig06]. Similar optimization problems under hard constraints have been proposed and efficiently addressed using the adjoint method [Perez2015rod, Zehnder2016curve, Perez2017surfaces, Ly2018shell, Malomo2018flexmaps, Panetta2019]. While these adjoint-based approaches are efficient, their implementation tends to be significantly more complicated because the nonlinear force constraints need to be differentiated, thus involving the Hessian of the objectives of elastic materials. Besides complexity, prior work has highlighted additional challenges introduced by this Hessian, related to correctness, accuracy, robustness, and efficiency [Panetta2019, Shi2023curls]. As such, similar to the work of [Twigg:2011:OSS:2019406.2019437], we formulate our method without explicitly encoding the Hessian into the rest shape optimization. To avoid these complications for the inversions, Choi et al. [Choi2024] proposed using the finite difference method to approximate the gradient with respect to the curvatures while Liu [Liu2022gravity] treated a simulator as a black-box using search directions approximated with deformation gradients, although such schemes are known to be rather inefficient.
@@ 2.2.3 Global-Local Initialization
Hsu et al. [Hsu2022sag] proposed a global-local, two-stage initialization method that first solves a global linear system to find optimal forces and then locally adjusts parameters on the rest shape, material stiffness, and internal state such that these parameters become consistent with the computed forces. Compared to the work of Twigg and Kačić-Alesić [Twigg:2011:OSS:2019406.2019437] and our own, which need to solve a global nonlinear system iteratively, the two-stage initialization [Hsu2022sag] requires solving a global linear system only once which reduces the computational cost significantly. This two-stage initialization was also extended to support hair simulation [Hsu2023sag] with the elastic rod model of [Kugelstadt2016rod]. In practice, while the two-stage initialization is efficient, forces computed via the single global linear solve are not necessarily feasible with only the modified rest shape parameters of local elements. As such, one would typically need to compromise with non-zero forces failing to achieve static equilibrium or locally modified material stiffness parameters deviating from the user-desired strand dynamics, as discussed in [Hsu2022sag, Hsu2023sag].
@@
In addition, it is nontrivial to apply the global-local initialization approach [Hsu2022sag, Hsu2023sag] to DER because its bending and twisting formulations do not satisfy the requirements of the global-local initialization. First, the bending and twisting formulations are tightly coupled in DER, and thus the local (minimal) stencil for bending does not necessarily preserve the angular momentum by itself (we confirmed this with the optimal forces generated by our method). This fact makes it difficult to define force elements, which are used to form the global linear system. Furthermore, assuming the use of the local stencil as a force element, the bending/twisting couples the rest length parameters with neighboring ones, transforming the (ostensibly) local steps into a global problem.
2.2.4 Differences from [Twigg:2011:OSS:2019406.2019437]
@@
Twigg and Kačić-Alesić [Twigg:2011:OSS:2019406.2019437] proposed a rest shape optimization framework for sag-free mass-spring systems. Their method formulates an objective based on the $L^{2}$ norm of forces with a related regularizer, and optimizes the rest length and rest angle under box constraints for the mass-spring networks. While this work has a goal similar to ours, there are two key differences. First, instead of mass-spring systems, we use DER (which has been validated through physical experiments to generate dynamics in accordance with theory [Bergou2008, Bergou:2010:DVT:1778765.1778853, jawed2018primer, Romero2021validation]) and therefore we optimize the rest shape parameters of the DER objectives: rest length, rest curvature, and rest twist. In addition, we also impose corresponding box constraints (based on our analysis and numerical experiments) on the rest shape parameters to prevent significant changes in their values and thus avoid introducing stability issues. Second, instead of the $L^{2}$ norm of forces, we formulate an objective based on kinetic energy. While these are conceptually similar, our kinetic-energy-based objective can correctly reflect the mass/inertia of materials and consistently evaluate the energy contributions of each component in the system. Moreover, our kinetic-energy-based formulation has much better numerical conditioning than the force-norm-based one (see Sec. 4.1.5), making it possible to use double-precision floating-point when solving the optimization problem. By contrast, the force-norm-based objective can be numerically ill-conditioned such that, to avoid convergence failure during the optimization, extended-precision is required (such as the quad-precision adopted in [Twigg:2011:OSS:2019406.2019437]). These features are not standard nor officially supported in C++20, for example, and significantly slow down the computations.
@@ 3 Discrete Elastic Rods Preliminaries
In preparation for the rest shape optimization approach we develop in Sec. 4, we briefly review the energy objectives in the DER formulation and their gradients, which play key roles in forward simulation [Bergou2008, Bergou:2010:DVT:1778765.1778853]. Here, as our focus is on the static equilibrium case, we do not consider plastic deformations or tearing, and cover details only directly relevant to our rest shape optimizer. We refer readers to the book on DER [jawed2018primer] for the underlying principles including space/time-parallel transport and to the previous work for derivation and details of the gradient and Hessian [Panetta2019, Fei2019]. In addition, because we consider scenarios without any contacts among elastic rods and other objects, each rod can independently be processed in both forward simulation and rest shape optimization. As such, we describe formulations for a single elastic rod below.
@@
We consider an elastic rod discretized with $N$ vertices (which are sufficiently well distributed without forming acute angles) whose stacked positions are denoted by $\mathbf{x}=[\mathbf{x}_{0}^{T},\ldots,\mathbf{x}_{N-1}^{T}]^{T}\in\mathbb{R}^{3N}$ (each vertex position is three-dimensional) and $(N-1)$ connecting edges whose stacked edge angles are $\bm{\theta}=[\bm{\theta}_{0},\ldots,\bm{\theta}_{N-2}]^{T}\in\mathbb{R}^{N-1}$ (each edge angle is one-dimensional). The generalized positions of the rod can be defined by interleaving the vertex positions and edge angles as $\mathbf{q}=[\mathbf{x}_{0}^{T},\bm{\theta}_{0},\ldots,\bm{\theta}_{N-2},\mathbf{x}_{N-1}^{T}]^{T}\in\mathbb{R}^{4N-1}$. This specific variable arrangement of $\mathbf{q}$ leads to a banded Hessian for the DER objectives with time-parallel transport (which also accommodates trivial parallel computations [Kaldor2010]) and thus enables efficient implicit integration [Bergou:2010:DVT:1778765.1778853], via Cholesky/LDLT-based direct linear solves [Huang2023hair]. Along with the positional variables, we also define stacked velocities for the vertices as $\mathbf{\dot{x}}=[\mathbf{\dot{x}}_{0}^{T},\ldots,\mathbf{\dot{x}}_{N-1}^{T}]^{T}\in\mathbb{R}^{3N}$, stacked angular velocities for the edges $\bm{\dot{\theta}}=[\bm{\dot{\theta}}_{0},\ldots,\bm{\dot{\theta}}_{N-2}]^{T}\in\mathbb{R}^{N-1}$, and generalized velocities $\mathbf{\dot{q}}=[\mathbf{\dot{x}}_{0}^{T},\bm{\dot{\theta}}_{0},\ldots,\bm{\dot{\theta}}_{N-2},\mathbf{\dot{x}}_{N-1}^{T}]^{T}\in\mathbb{R}^{4N-1}$. For simplicity, we assume perfectly circular cross-sections with a constant radius, constant density, and constant material stiffness over the strand. In addition, we initialize reference frames on each edge via space-parallel transport propagated from a randomly chosen reference frame on the first edge (i.e., any reference frame on the first edge, as long as it forms the $SO(3)$ basis with the unit tangent vector, gives identical results for the dynamics) and then compute material frames based on the reference frames and edge angles $\bm{\theta}$ [Bergou2008, Bergou:2010:DVT:1778765.1778853, jawed2018primer]. While, in the previous work [Bergou2008, Bergou:2010:DVT:1778765.1778853, jawed2018primer], subscript and superscript indices are used for the variables defined on vertices and edges, respectively, we use only subscript indices to avoid confusion with common superscripts (e.g., exponentiation, inverse, transpose, and time index), and because we can identify whether variables are defined on a vertex or an edge from their definition.
@@
While the dimensionality of $\mathbf{q}$ is $(4N-1)$ in 3D, we focus on an elastic rod with its root end minimally clamped as this is one of the simplest and most common settings, e.g., for hairs attached to a head (although it would be possible to support both ends clamped [Hafner2023rod]). To this end, we fix $\mathbf{x}_{0},\bm{\theta}_{0}$, and $\mathbf{x}_{1}$ or update them in a prescribed way (below, we use the term “fix” regardless of whether these variables are fixed in place or moved in a prescribed way) so that the elastic rod has an anchor enabling stretching/bending/twisting to work meaningfully, i.e., these variables work as a Dirichlet boundary condition to eliminate DOFs for pure rigid motions. This treatment corresponds to eliminating the DOFs for $\mathbf{x}_{0},\bm{\theta}_{0}$, and $\mathbf{x}_{1}$, and thus under our minimal root-end clamping the resulting number of free DOFs is $(4N-8)=(4N-1)-(3+1+3)$.
@@
A single forward simulation step of DER can be written as an energy minimization problem [Martin:2011:EEM:2010324.1964967, Gast2015]:
@@
where the objective of DER, $E_{\mathrm{DER}}(\mathbf{q})$ consists of the objective terms for inertia $E_{\mathrm{inertia}}(\mathbf{q})$, stretch $E_{\mathrm{stretch}}(\mathbf{x})$, bend $E_{\mathrm{bend}}(\mathbf{q})$, and twist $E_{\mathrm{twist}}(\mathbf{q})$, and can be defined as

### v2

@@ 1 Introduction
From human hairs to cables for electronic devices, elastic strands can be found everywhere in our daily lives, and modeling such thread-like materials is of critical importance in various domains, including computer graphics, mechanical engineering, and medical applications. The Discrete Elastic Rods (DER) formulation [BWR08, BAV10] is one of the most popular approaches to accurately and efficiently simulate such elastic strands and has demonstrated its effectiveness in a wide range of applications involving both forward simulations [CAR09, DKH10, ACB13, KTS14, SJM17, FBGZ19, LSD22, Dav23] and inverse problems [PTC15, ZCT16, POT17, SMGT18, MPI18, XKCB18, PKLI19, NPTB22, RKP22, CZB24].
When modeling and designing elastic strands, a well-known problem is that such strands will immediately sag due to gravity when the simulation starts [TKA11, HWP23], which ruins the specific strand shapes (e.g., hairstyles) that were either carefully designed by users, technical artists, and designers [IBB19, Liu22] or captured from real human hair or wigs [LLR13]. This issue occurs because the simplest way to model the strands is to focus solely on the geometry, neglecting the role of forces or dynamics; that is, the strand’s initial shape is treated as its rest shape [KE22]. While one can largely mitigate this sagging effect by significantly (and artificially) increasing the material stiffness, this approach makes the elastic strands excessively stiff and sacrifices their natural dynamic response, thus negatively impacting the visual result.
To address this sagging problem in the DER framework, we propose a rest shape optimization approach that optimizes the rest shape parameters (rest length, rest curvature, and rest twist) of the input strands such that the input shape is preserved at static equilibrium under simulation (i.e., the strand shape deformed due to simulation coincides with the input shape). To find optimal rest shape parameters, we formulate a minimization problem based on the kinetic energy. Given the greater number of degrees of freedom (DOFs) for the rest shape parameters compared to the DOFs of DER strands, we incorporate a regularizer to avoid creating underdetermined systems. In addition, we impose box constraints on the rest shape parameters to prevent significant changes from their original values, ensuring the stability of the system. We solve the resulting constrained nonlinear minimization problem via the Gauss-Newton (GN) algorithm (and thus without the Hessian of the DER objectives) augmented with penalty methods and line search. Fig. Rest Shape Optimization for Sag-Free Discrete Elastic Rods demonstrates the efficacy of our rest shape optimization.
@@ 2.1 Elastic Rod Simulation
Due to the one-dimensional structure of thin elastic strands, it is typically difficult to accurately simulate these materials with general finite element methods (FEM) designed for volumetric structures [KE22], and specialized approaches have been investigated. Given that common strand-like materials exhibit higher stiffness in stretching modes compared to bending and twisting, early work assumed inextensibility via e.g., constraints based on Lagrange multipliers [GHF07] with linear-time factorization and triangular solves [Bar96], a tridiagonal matrix formulation [HH13], and geometric stiffness [TNGF15, ATK17], reduced multi-body solvers [Had06, Fea16], and position-based dynamics (PBD) [MHHR07, BMM17]. To support the characteristic dynamics of elastic strands, the traditional mass-spring framework [LBOK13] was extended with altitude springs [SLF08], bending springs [IMP13], torsion energy [SWP23], and ghost rest shapes [HZS24]. To simulate one-dimensional strands along with other deformables in a unified way, supporting torsional effects is critical, and therefore discretization elements that can capture twists (e.g., oriented particles [MC11] and elastons [MKB10]) have also been presented. For highly constrained strand-like systems, further specialized approaches have been proposed with, e.g., Eulerian nodes [SJLP11, SBRBO20], long range constraints [MCMJ17], and unilateral distance constraints [MCJM18].
Since the introduction of the Cosserat theory into graphics [Pai02], it has been adapted to dynamical systems of strands [ST07] and network structures [ST09], combined with PBD using ghost particles [USS14], and further extended to support edge rotations via quaternions in the PBD framework [KS16]. The work of Kugelstadt and Schömer [KS16] has been extended to remove the ambiguity in the Darboux vector [HWP23], to handle stiff problems via compliant constraints while utilizing the tree structures [DKWB18], to preserve the volume of elastic strands [ARM19], and to achieve consistent dynamics [SMSH18] using projective dynamics [BML14]. Elastic rod simulation based on the Cosserat theory [ST07] has also been extended to support rh-adaptivity [WCU20] and a compact representation of the DOFs [ZLW22]. A curvature-based reduced representation has also been presented by Bertails et al. [BAC06] and extended to generate smoother dynamics via clothoids [CBD13].
While various approaches have been presented for simulating elastic strands, DER [BWR08, BAV10], which discretizes an elastic strand with vertex positions and edge angles in maximal coordinates, is one of the most popular approaches due to its simplicity, efficiency, and flexibility, as employed in a wide range of applications. As such, we use DER as our elastic strand simulator.
@@ 2.2.1 Nonlinear Force Solve
Solving Newton’s second law of motion for (typically nonlinear) forces and zero acceleration is equivalent to achieving zero net force and thus static equilibrium. Hadap [Had06] proposed using inverse dynamics [Fea16] to solve the nonlinear equation on the forces and thus enable sag-free simulation. While inverse dynamics works for reduced multibody systems, in general, it is not clear how to apply it for general simulation systems represented in maximal coordinates (e.g., FEM-based deformables and DER).
The nonlinear equations on the forces can be solved, e.g., with Newton-type optimizers [WWY15, MWW18], although Iben [IBB19] aimed to cancel gravity using spring forces via local analytical solutions within the mass-spring systems [IMP13]. However, in practice, it is necessary to consider a global system to fully cancel gravity. A similar technique was also employed for sag-free mass-spring systems by introducing additional artificial springs (which can potentially introduce unnatural forces) [HZS24].
Chen et al. [CZXZ14] employed an Asymptotic Numerical Method (ANM), which is a homotopy approach to solving nonlinear equations, for inverse elastic shape design with incompressible neo-Hookean materials. They demonstrated orders of magnitude faster convergence than one of the Newton-type optimizers, Levenberg–Marquardt algorithm (LMA) [NW06]. Later, Jia [Jia21] extended ANM to support more general material types and also demonstrated its efficiency for forward simulation compared to optimization-based integrators [MTGG11, GSS15] combined with Newton’s method [NW06] (although ANM can yield solutions that are local maxima, unlike those generated by the optimization-based integrators). ANM was also employed to investigate the equilibrium of elastic rods [LMR13]. While ANM can be more efficient than the Newton-type optimizers [NW06], it is not clear how to support box constraints, which are indispensable to constrain rest shape parameters within physically valid ranges and avoid introducing stability issues.
While canceling gravity forces in DER has been previously attempted by solving force equations or equivalent minimization [AX15, LSD22], their approach lacked algorithmic and formulation specifics, such as exact objectives and DER force Jacobians with respect to the rest shape parameters. Additionally, their approach did not support box constraints which are critical to avoid significant rest shape changes and thus ensure stable simulations.
@@ 2.2.2 Constrained Minimization
Derouet-Jourdan et al. [DJBDT10] presented a method for achieving static equilibrium of elastic strands, simulated with the curvature-based discretization [BAC06], in 2D. They formulated an objective consisting of curvatures and material parameters and then optimized these parameters via linear solves on the curvature and stiffening/lightening of materials. Derouet-Jourdan et al. [DJBDDT13] also achieved static equilibrium of curvature-based hair strands [BAC06] in 3D under frictional contacts. Along the same lines, inverse approaches for curvature-based elastic strands have been further investigated [BDDJRL18, CBDNR20, HB21, HB23].
Twigg and Kačić-Alesić proposed optimizing rest shape parameters via force-norm minimization to achieve static equilibrium of mass-spring systems [TKA11]. Our work is most closely related to theirs, so we clarify the differences in detail in Sec. 2.2.4. Skouras et al. [STBG12] presented an optimization approach that treats the nonlinear force equations as hard constraints and solves the optimization via the augmented Lagrangian method [NW06]. Similar optimization problems under hard constraints have been proposed and efficiently addressed using the adjoint method [PTC15, ZCT16, POT17, LCBD18, MPI18, PKLI19]. While these adjoint-based approaches are efficient and can ensure local minima (not maxima), their implementation tends to be significantly more complicated because the nonlinear force constraints need to be differentiated, thus involving the Hessian of the objectives of elastic materials. Besides complexity, prior work has highlighted additional challenges introduced by this Hessian, related to correctness, accuracy, robustness, and efficiency [PKLI19, SWP23]. As such, similar to the work of [TKA11], we formulate our method without explicitly encoding the Hessian into the rest shape optimization. To avoid these complications for the inversions, Choi et al. [CJSJ24] proposed using the finite difference method to approximate the gradient with respect to the curvatures while Liu [Liu22] treated a simulator as a black-box using search directions approximated with deformation gradients, although such schemes are known to be rather inefficient.
@@ 2.2.3 Global-Local Initialization
Hsu et al. [HTYW22] proposed a global-local, two-stage initialization method that first solves a global linear system to find optimal forces and then locally adjusts parameters on the rest shape, material stiffness, and internal state such that these parameters become consistent with the computed forces. Compared to the work of Twigg and Kačić-Alesić [TKA11] and our own, which need to solve a global nonlinear system iteratively, the two-stage initialization [HTYW22] requires solving a global linear system only once which reduces the computational cost significantly. This two-stage initialization was also extended to support hair simulation [HWP23] with the elastic rod model of [KS16]. In practice, while the two-stage initialization is efficient, forces computed via the single global linear solve are not necessarily feasible with only the modified rest shape parameters of local elements. As such, one would typically need to compromise with non-zero forces failing to achieve static equilibrium or locally modified material stiffness parameters deviating from the user-desired strand dynamics, as discussed in [HTYW22, HWP23].
@@
In addition, it is nontrivial to apply the global-local initialization approach [HTYW22, HWP23] to DER because its bending and twisting formulations do not satisfy the requirements of the global-local initialization. First, the bending and twisting formulations are tightly coupled in DER, and thus the local (minimal) stencil for bending does not necessarily preserve the angular momentum by itself (we confirmed this with the optimal forces generated by our method). This fact makes it difficult to define force elements, which are used to form the global linear system. Furthermore, assuming the use of the local stencil as a force element, the bending/twisting couples the rest length parameters with neighboring ones, transforming the (ostensibly) local steps into a global problem. As such, it is necessary to modify the global-local initialization to apply it to DER although such an approach would fail to achieve static equilibrium (see Sec. 5.6).
2.2.4 Differences from [TKA11]
@@
Twigg and Kačić-Alesić [TKA11] proposed a rest shape optimization framework for sag-free mass-spring systems. Their method formulates an objective based on the $L^{2}$ norm of forces with a related regularizer, and optimizes the rest length and rest angle under box constraints for the mass-spring networks. While this work has a goal similar to ours, there are two key differences. First, instead of mass-spring systems, we use DER (which has been validated through physical experiments to generate dynamics in accordance with theory [BWR08, BAV10, JNO18, RLR21]) and therefore we optimize the rest shape parameters of the DER objectives: rest length, rest curvature, and rest twist. In addition, we also impose corresponding box constraints (based on our analysis and numerical experiments) on the rest shape parameters to prevent significant changes in their values and thus avoid introducing stability issues. Second, instead of the $L^{2}$ norm of forces, we formulate an objective based on kinetic energy. While these are conceptually similar, our kinetic-energy-based objective can correctly reflect the mass/inertia of materials and consistently evaluate the energy contributions of each component in the system. Moreover, our kinetic-energy-based formulation has much better numerical conditioning than the force-norm-based one due to the proper consideration of the energy contributions (see Sec. 4.6), making it possible to use double-precision floating-point when solving the optimization problem. By contrast, the force-norm-based objective can be numerically ill-conditioned such that, to achieve sufficiently small error during the optimization, extended-precision would be required (such as the quad-precision adopted in [TKA11]). These features are not standard nor officially supported in C++20, for example, and significantly slow down the computations.
@@ 3 Discrete Elastic Rods Preliminaries
In preparation for the rest shape optimization approach we develop in Sec. 4, we briefly review the energy objectives in the DER formulation and their gradients, which play key roles in forward simulation [BWR08, BAV10]. Here, as our focus is on the static equilibrium case, we do not consider contacts, plastic deformations, or tearing, and cover details only directly relevant to our rest shape optimizer. We refer readers to the book on DER [JNO18] for the underlying principles including space/time-parallel transport and to the previous work for derivation and details of the gradient and Hessian [PKLI19, FBGZ19]. In addition, because we consider scenarios without any contacts among elastic rods and other objects, each rod can independently be processed in both forward simulation and rest shape optimization. As such, we describe formulations for a single elastic rod below.
@@
We consider an elastic rod discretized with $N$ vertices (which are sufficiently well distributed without forming acute angles) whose stacked positions are denoted by $\mathbf{x}=[\mathbf{x}_{0}^{T},\ldots,\mathbf{x}_{N-1}^{T}]^{T}\in\mathbb{R}^{3N}$ (each vertex position is three-dimensional) and $(N-1)$ connecting edges whose stacked edge angles are $\bm{\theta}=[\bm{\theta}_{0},\ldots,\bm{\theta}_{N-2}]^{T}\in\mathbb{R}^{N-1}$ (each edge angle is one-dimensional). The generalized positions of the rod can be defined by interleaving the vertex positions and edge angles as $\mathbf{q}=[\mathbf{x}_{0}^{T},\bm{\theta}_{0},\ldots,\bm{\theta}_{N-2},\mathbf{x}_{N-1}^{T}]^{T}\in\mathbb{R}^{4N-1}$. This specific variable arrangement of $\mathbf{q}$ leads to a banded Hessian for the DER objectives with time-parallel transport (which also accommodates trivial parallel computations [KJM10]) and thus enables efficient implicit integration [BAV10], via Cholesky/LDLT-based direct linear solves [HYW23]. Along with the positional variables, we also define stacked velocities for the vertices as $\mathbf{\dot{x}}=[\mathbf{\dot{x}}_{0}^{T},\ldots,\mathbf{\dot{x}}_{N-1}^{T}]^{T}\in\mathbb{R}^{3N}$, stacked angular velocities for the edges $\bm{\dot{\theta}}=[\bm{\dot{\theta}}_{0},\ldots,\bm{\dot{\theta}}_{N-2}]^{T}\in\mathbb{R}^{N-1}$, and generalized velocities $\mathbf{\dot{q}}=[\mathbf{\dot{x}}_{0}^{T},\bm{\dot{\theta}}_{0},\ldots,\bm{\dot{\theta}}_{N-2},\mathbf{\dot{x}}_{N-1}^{T}]^{T}\in\mathbb{R}^{4N-1}$. For simplicity, we assume perfectly circular cross-sections with a constant radius, constant density, and constant material stiffness over the strand. In addition, we initialize reference frames on each edge via space-parallel transport propagated from a randomly chosen reference frame on the first edge (i.e., any reference frame on the first edge, as long as it forms the $SO(3)$ basis with the unit tangent vector, gives identical results for the dynamics) and then compute material frames based on the reference frames and edge angles $\bm{\theta}$ [BWR08, BAV10, JNO18]. While, in the previous work [BWR08, BAV10, JNO18], subscript and superscript indices are used for the variables defined on vertices and edges, respectively, we use only subscript indices to avoid confusion with common superscripts (e.g., exponentiation, inverse, transpose, and time index), and because we can identify whether variables are defined on a vertex or an edge from their definition.
@@
While the dimensionality of $\mathbf{q}$ is $(4N-1)$ in 3D, we focus on an elastic rod with its root end minimally clamped as this is one of the simplest and most common settings, e.g., for hairs attached to a head (although it would be possible to support both ends clamped [HB23]). To this end, we fix $\mathbf{x}_{0},\bm{\theta}_{0}$, and $\mathbf{x}_{1}$ or update them in a prescribed way (below, we use the term “fix” regardless of whether these variables are fixed in place or moved in a prescribed way) so that the elastic rod has an anchor enabling stretching/bending/twisting to work meaningfully, i.e., these variables work as a Dirichlet boundary condition to eliminate DOFs for pure rigid motions. This treatment corresponds to eliminating the DOFs for $\mathbf{x}_{0},\bm{\theta}_{0}$, and $\mathbf{x}_{1}$, and thus under our minimal root-end clamping the resulting number of free DOFs is $(4N-8)=(4N-1)-(3+1+3)$.
@@
A single forward simulation step of DER can be written as an energy minimization problem [MTGG11, GSS15]:
@@
where the objective of DER, $E_{\mathrm{DER}}(\mathbf{q})$ consists of the objective terms for inertia $E_{\mathrm{inertia}}(\mathbf{q})$, stretching $E_{\mathrm{stretch}}(\mathbf{x})$, bending $E_{\mathrm{bend}}(\mathbf{q})$, and twisting $E_{\mathrm{twist}}(\mathbf{q})$, and can be defined as